OpenStudy (anonymous):

does someone know the answer to (-2)4/3

OpenStudy (lukecrayonz):

Yes, follow the rules of pemdas, -8/3

OpenStudy (anonymous):

(-2)4/3=-8/3=-2.66666....

OpenStudy (anonymous):

thanks i knew i was correct

OpenStudy (anonymous):

i need assistance on (81)1/4 x (2)-1(exponent)